OpenStudy (anonymous):

You are fencing a rectangular garden at your house. The length of the rectangle is 3 more than the width. The perimeter of the rectangular garden can be no more than 40 feet. What are the dimensions of the rectangular garden? Solve with inequality

OpenStudy (anonymous):

Hi...

OpenStudy (anonymous):

Any work on this so far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can we suppose L be length and W be width of the rectangle ?

OpenStudy (anonymous):

We have given L = W + 3. right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now we have perimeter = L + W. right ? so we get (W + 3) + W = 40. right ?

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

Now adding like terms we get 2W + 3 = 40. right ?

OpenStudy (anonymous):

ok but what would be the inequality?

OpenStudy (anonymous):

Inequality is for length is 3 more than width this means L > W + 3. got it now ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So by adding -3 on both side we get 2W = 37. right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dividing both side by 2 we get W = 37/2. okay ?

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

So your width should be 37/2. Okay ?

OpenStudy (anonymous):

Now plug W = 37/2 in L = W + 3 to get the value of L. ok??

OpenStudy (anonymous):

L = 37/2 + 3 = 43/2. Okay ?

OpenStudy (anonymous):

So your dimentions must be length is 43/2 feet and width is 37/2 feet. Okay ?

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

you are most Welcome :) Have a nice day ahead...!

OpenStudy (anonymous):

okay you too

OpenStudy (anonymous):

Good bye :)

OpenStudy (anonymous):

byee